# Self-assembled filament layers in drying sessile droplets: from morphology to electrical conductivity

### v1

@@ I Introduction
Drying droplets are a versatile approach for assembling functional materials - such as nanowires, carbon nanotubes, and conducting polymers - into organized structures for a wide range of emerging applications. These include energy devices, sensors, neuromorphic circuits, biomedical systems, and printed electronics [Steinberger2024, Rao2022, Kwon2020, Wang2021]. In this process, functional nanomaterials are dispersed in liquid ink droplets and deposited onto a substrate via techniques like inkjet, aerosol-jet, or spray printing. As the solvent evaporates, internal convective flows, interfacial forces, and particle interactions drive the dispersed materials to self-organize into patterned, anisotropic, or hierarchical arrangements [Park2024, Dugyala2015]. Such evaporation-induced self-assembly enables the formation of tailored micro- and nanostructured films that are attractive for advanced manufacturing.
@@
Device performance is strongly determined by the morphology of the deposited layer. In printed electronics, evaporation-induced self-assembly enables the fabrication of flexible and cost-effective filament layers [Kamyshny2019, Li2022]. For filament-based materials such as carbon nanotubes, silver nanowires, or conducting polymers, alignment reduces the number of filament crossings and enhances in-plane charge transport, while the quality of contact points and phase organization further improve conductivity and mechanical stability [Ren2012, Xu2020]. In addition to alignment, filament distribution and surface roughness critically affect electronic performance: rough or coiled deposits can create leakage pathways and limit efficiency, whereas embedding or planarization techniques reduce roughness and improve device reliability [HosseinzadehKhaligh2014].
@@
The final dried morphology of the deposit layer results from a complex interplay of fluid dynamics, interfacial effects, and particle interactions. The flows inside the droplets are mainly dependent on the evaporation modes (contact line dynamics) and evaporation regimes (evaporation rate/flux distribution). When a sessile droplet dries, the contact line typically follows one of three different modes: the constant contact angle mode (CCA, mobile contact line), the constant contact radius mode (CCR, pinned contact line), or a combination of these two, known as the stick-slide mode. In practice, printed droplets often exhibit stick-slide behavior, characterized by an initial period of pinning followed by depinning. Evaporation regimes are generally classified as diffusion-limited or reaction-limited. Under diffusion-limited conditions, vapor transport away from the interface controls the evaporation rate. At low contact angles, the evaporation flux diverges near the contact line, generating strong outward capillary flows that promote ring-like deposits [Deegan1997]. By contrast, in reaction-limited (interface-limited) drying, interfacial kinetics govern the evaporation rate. The evaporation flux distribution along the droplet surface is more uniform, yielding weaker capillary flows and a more homogeneous deposition [Wilson2023, Larsson2023]. The morphology of the dried layer is also highly sensitive to parameters such as solvent composition or ambient conditions [Hu2006], particle concentration [Baldwin2011], aspect ratio [Dugyala2015] and stiffness [Park2024]. Depending on these factors, deposits can range from uniform coatings or aligned networks to uneven ring stains caused by the coffee-ring effect (CRE), where particles accumulate at the contact line. Uniform deposits are often desirable, especially in coating processes and printing applications [Steinberger2024].
@@
The electrical conductivity of the deposited layer is one of the key properties for applications in printed electronics [Bi2024]. Percolation theory plays a key role in understanding and optimizing the conductivity of such morphologies, describing the transition from isolated clusters to a continuous, conductive network, governed by filament concentration, shape, and distribution [Berg2024]. Filaments are particularly effective in achieving percolation at low concentrations, as their elongated shapes increase the likelihood of overlap and network formation. Their anisotropic geometry also promotes orientational ordering during deposition, which can enhance electrical, optical, and mechanical properties. This makes them attractive for applications requiring high conductivity and material efficiency [Cao2020].
@@
Experimental and theoretical approaches provide valuable insights, but face limitations in isolating and controlling the parameters of drying filament droplets. Numerical simulations overcome these constraints by enabling the systematic variation of individual parameters while keeping others fixed. While spherical particles are well-studied [Xie2018, Xie2025, Hamjah2021], computational modeling of fluid-coupled rod-like structures forming deposition patterns during drying droplets remains largely unexplored [Kravchenko2020, Kolegov2025]. To date, simulations of such systems have often been simplified, for instance by using Monte Carlo (MC) methods that treat filaments as random stick networks without incorporating the fluid dynamics of the evaporating droplet [Hicks2018, Fata2020].
@@
In this work, we couple the lattice Boltzmann method for fluid dynamics with a bead-spring approach for filaments. Our approach enables the simulation of fluid flow, particle interactions, and network formation, providing insights into filament alignment, clustering, and percolation behavior following solvent drying. We consider the filament suspension in a dilute limit, where self-pinning is negligible, and focus on the stick–slide evaporation mode: droplets evaporate in CCR mode until a critical receding angle $\theta_{re}$ is reached, after which they transition to CCA mode [Orejon2011, Wilson2023]. We investigate how filament length, concentration, stiffness, and evaporation conditions affect percolative network formation through a systematic parameter study. We then analyze the nematic order, radial density profiles, percolation probability, and electrical conductivity of the deposit.
@@ II.1 Color-gradient lattice Boltzmann model
The fluid solver is based on the three-dimensional lattice Boltzmann method with $19$ discrete velocities (D3Q19) [Krger2017, Benzi1992]. We describe the fluid forming the droplet and the surrounding fluid as independent components. Each component $k\in\{1,2\}$ is described by a distribution function $f_{i}^{k}(\mathbf{x},t)$, which evolves according to the lattice Boltzmann equation,
@@
where $c_{s}=\frac{1}{\sqrt{3}}\frac{\Delta x}{\Delta t}$ is the speed of sound, $\Omega_{i}^{k}$ is the collision operator, $\mathbf{c}_{i}$ are the discrete velocity vectors, and $i=1,\dots,19$. We set the timestep $\Delta t$ and the lattice constant $\Delta x$ to unity for simplicity. The macroscopic density and velocity of component $k$ are given by
@@
with $\rho_{0}$ a reference density.
The CGLB method, enables the simulation of immiscible or partially miscible fluids by modeling interfacial tension and enforcing phase segregation through three substeps: relaxation, perturbation, and recoloring [Gunstensen1991, Leclaire2017, Aursj2015]. The total collision operator $\Omega_{i}$ acting on the total distribution function $f_{i}$ is decomposed as
@@
Here, $\tau$ is the relaxation time and the kinematic viscosity is $\nu=c_{s}^{2}(\tau-0.5)$. The equilibrium distribution is given by
@@
where $\rho=\rho_{1}+\rho_{2}$ and $\mathbf{u}$ is the mixture velocity. The lattice weights $w_{i}$ for D3Q19 are
@@
Perturbation: Interfacial tension is introduced via a perturbation proportional to the color gradient $\mathbf{G}$:
@@
where $\phi_{i}$ is the angle between $\mathbf{G}$ and $\mathbf{c}_{i}$. The parameter $A$ sets the surface tension magnitude $\sigma$ according to $A=\frac{9\sigma}{4\tau}$ [Leclaire2017], while the lattice-dependent weights $B_{i}$ are defined as
@@
To describe wetting on the substrate, we follow the work of Akai et al. [Akai2018]. We dynamically adjust the fluid interface normal $\mathbf{n}_{f}$ at the triple line to enforce the contact angle condition. The algorithm computes two candidate normals $\mathbf{v}_{0}$ and $\mathbf{v}_{1}$ based on the geometric relationship between the fluid interface normal $\mathbf{n}_{f}$ and the wall normal $\mathbf{n}_{w}$:
@@ II.2 Evaporation model
Fluid evaporation is implemented following the recent work of Nath et al. [Nath2025]. This reaction-limited evaporation model simulates phase change by converting distribution populations at the liquid-vapor interface. This approach is adaptable to any evaporation regime with a known analytical evaporation flux expression and is computationally efficient, as calculations are confined to the interface. Evaporation is modeled by converting resting populations at the interface sites of unit volume $\mathbf{x}\in\mathbf{x}_{I}$:
$f_{0}^{r,b}(\mathbf{x},t)^{\text{new}}=f_{0}^{r,b}(\mathbf{x},t)\pm\varphi\Delta t$ (15)
The surface mass flux $\Xi$ across each site is linked to a volumetric sink rate $\varphi=\frac{d\rho}{dt}=S\Xi$, with a correction factor $S:=1/3$ accounting for discretization errors. During each timestep, mass is subtracted from one fluid component and added to another, preserving total mass and momentum.
In printing technologies, evaporation typically occurs in an intermediate regime between the diffusion-limited and reaction-limited limits, which can be adjusted by masking the droplet, varying the relative humidity or the ambient pressure or by imposing an airflow [Majewski2025].
In the reaction-limited regime, interfacial kinetics - rather than vapor diffusion - sets the evaporation rate. This commonly occurs on heated substrates or at low pressure [MURISIC2011, Nath2025]. The surface mass flux $\Xi$ follows the Hertz–Knudsen (HK) relation. Using the Clausius-Clapeyron law, the HK relation is given by
$\Xi=\frac{\Lambda\,\rho_{v}\,\mathcal{L}}{T_{\text{sat}}^{3/2}}\sqrt{\frac{M_{v}}{2\pi R_{g}}}\,\left[T_{i}-T_{\text{sat}}\right],$ (16)
with $\Lambda$ the accommodation coefficient, $\rho_{v}$ the saturation vapor density, $M_{v}$ the molecular mass of the vapor, $R_{g}$ the universal gas constant, $\mathcal{L}$ the latent heat of vaporization, $T_{i}$ the liquid–vapor interface temperature, and $T_{\text{sat}}$ the saturation temperature. In the non-equilibrium one-sided model (NEOS) [Larsson2023, MURISIC2011], under the thin-droplet assumption, with negligible substrate thickness and curvature effects, the evaporation flux $J$ for thin droplets reduces to
@@
In the case of diffusion-limited evaporation of a sessile droplet, vapor transport is assumed quasi-steady and purely diffusive, governed by the Laplace equation [Deegan1997, Hu2002]
@@
Here, $D_{v}$ is the vapor diffusivity in air and $u(\mathbf{x})$ the vapor concentration field. The boundary conditions are $u=u_{s}$ (saturated vapor concentration) at the droplet surface and $u=u_{\infty}$ (ambient vapor concentration) far away. The droplet is modeled as a spherical cap (small Bond number), and viscous forces are negligible (low capillary number). The resulting evaporation flux $J(r)$ exhibits a divergence near the contact line. Nevertheless, the total evaporation rate remains finite because the singularity is integrable; in addition, real droplets often possess a thin precursor film that regularizes the divergence [Saxton2016, OBrien2024]. For small contact angles $(\theta\to 0)$, the evaporation flux is written as
@@
where $a$ is the droplet base radius and $r$ the radial position (see Fig. 1). For arbitrary angles, an approximate form is [Hu2002]
@@
in which $\theta(t)=2\arctan\left(\frac{h_{0}(t)}{a}\right)$ for a spherical droplet in case of semi-steady evaporation. Furthermore, the evaporation flux can be approximated as [MURISIC2011]
@@
In the diffusion-limited regime, the vapor fields of neighboring droplets overlap, leading to a shielding effect that reduces the local evaporation rate. The resulting local evaporative flux from the $j$-th droplet is expressed as [Wray2021]
$J_{j}(r,\xi)=J(r)\left[1-\frac{M\sqrt{b^{2}-a^{2}}}{2\pi\left(r^{2}+b^{2}-2rb\cos\xi\right)}\right],$ (22)
where $r$ is the radial coordinate from the center of the $k$-th droplet, $\xi$ the azimuthal angle measured from the axis connecting the droplet centers, and $F$ a shielding factor that quantifies the reduction of the evaporation flux due to the presence of a neighboring droplet. For two identical droplets of base radius $a$ separated by a center-to-center distance $b$, the shielding factor $F$ takes the analytical form [Wray2021]
@@ II.3 Filament model
The filaments within the droplet are modeled using a bead-spring representation. Each filament is constructed as a sequence of connected beads (monomers) linked by springs. Inter-monomer bonds are governed by the finite extensible nonlinear elastic (FENE) potential, known to accurately model real filament behavior [Li2012]:
@@
Here, $d$ is the separation between adjacent beads, $k_{s}=0.3$ is the spring constant, and $R_{m}=1.2$ is the maximum bond extension. To incorporate excluded volume effects, a Weeks-Chandler-Andersen (WCA) potential
@@
is added, where $\varepsilon=0.03$ is the depth of the potential well, and $d_{0}=1$ is the characteristic particle diameter (corresponding to the distance at which the potential vanishes). Stiffness of filaments is modeled using an approximate harmonic potential for second-nearest-neighbor interactions,
@@
Filaments are two-way coupled to the LB fluid via a local momentum-exchange scheme. Each bead experiences a hydrodynamic drag force [Ahlrichs1999, Jung2022]
@@
with $\gamma=6\pi\eta a_{0}$ representing the friction coefficient for a sphere of radius $a_{0}$ in a fluid of the dynamic viscosity $\eta$. Here, $\mathbf{v}_{\text{p}}$ denotes the bead velocity, while $\mathbf{v}_{\text{f}}$ is the fluid velocity interpolated at the bead’s position.
Conversely, the force exerted by the bead on the fluid is mapped back to the lattice using the method proposed by Ahlrichs and Dünweg [Ahlrichs1999]. This scheme distributes the momentum to the nearest lattice nodes via a regularized delta function. The update rule for the component distributions $f_{i}^{k}$ is given by
@@
The weight function $\delta(\mathbf{x},\mathbf{x}_{\text{p}})$ is the standard trilinear interpolation kernel with a stencil width of two lattice nodes in each spatial direction. This formulation locally injects momentum into the fluid while ensuring conservation.
The solvation force, introduced by Sega et al. [Sega2013], is used to confine filaments inside the droplet. It depends on the local gradients of the fluid-component densities,
@@
To capture capillary interactions, we calibrate the solvation magnitude $\kappa_{k}$ by balancing the net inward solvation force on the wetted segment $L_{w}$ against the capillary compression. Following the formulation established by Seong et al. [Seong2017], we approximate the theoretical capillary force $F_{c}$ as
@@
The prefactor $0.684$ is a geometric scaling factor adopted from [Seong2017].
The friction between filaments and the substrate critically influences the final deposition pattern during droplet evaporation. For adhesive surfaces, Derjaguin extended Amontons’ law by including an adhesion-related term [Derjaguin1934]. Accordingly, the friction force can be decomposed into external and internal contributions as [Gao2004]
@@
where $d$ is the distance from the substrate, and $\varepsilon=1.5\times 10^{-3}$ the potential well depth. This interaction restricts bead mobility, effectively generating internal friction. As evaporation proceeds, frictional forces can exceed solvation forces, immobilizing filaments and producing ring-like deposition patterns [Xie2018].

### v2

@@ I Introduction
Drying droplets are a versatile approach for assembling functional materials - such as nanowires, carbon nanotubes, and conducting polymers - into organized structures for a wide range of emerging applications. These include energy devices, sensors, neuromorphic circuits, biomedical systems, and printed electronics [63, 55, 36, 65]. In this process, functional nanomaterials are dispersed in liquid ink droplets and deposited onto a substrate via techniques like inkjet, aerosol-jet, or spray printing. As the solvent evaporates, internal convective flows, interfacial forces, and particle interactions drive the dispersed materials to self-organize into patterned, anisotropic, or hierarchical arrangements [53, 19]. Such evaporation-induced self-assembly enables the formation of tailored micro- and nanostructured films that are attractive for advanced manufacturing.
@@
Device performance is strongly determined by the morphology of the deposited layer. In printed electronics, evaporation-induced self-assembly enables the fabrication of flexible and cost-effective filament layers [30, 43]. For filament-based materials such as carbon nanotubes, silver nanowires, or conducting polymers, alignment reduces the number of filament crossings and enhances in-plane charge transport, while the quality of contact points and phase organization further improve conductivity and mechanical stability [56, 72]. In addition to alignment, filament distribution and surface roughness critically affect electronic performance: rough or coiled deposits can create leakage pathways and limit efficiency, whereas embedding or planarization techniques reduce roughness and improve device reliability [26].
@@
The final dried morphology of the deposit layer results from a complex interplay of fluid dynamics, interfacial effects, and particle interactions. The flows inside the droplets are mainly dependent on the evaporation modes (contact line dynamics) and evaporation regimes (evaporation rate/flux distribution). When a sessile droplet dries, the contact line typically follows one of three different modes: the constant contact angle mode (CCA, mobile contact line), the constant contact radius mode (CCR, pinned contact line), or a combination of these two, known as the stick-slide mode. In practice, printed droplets often exhibit stick-slide behavior, characterized by an initial period of pinning followed by depinning [67]. Evaporation regimes are generally classified as diffusion-limited or reaction-limited. Under diffusion-limited conditions, vapor transport away from the interface controls the evaporation rate. At low contact angles, the evaporation flux diverges near the contact line, generating strong outward capillary flows that promote ring-like deposits [15]. By contrast, in reaction-limited (interface-limited) drying, interfacial kinetics govern the evaporation rate. The evaporation flux distribution along the droplet surface is more uniform, yielding weaker capillary flows and a more homogeneous deposition [67, 37]. The morphology of the dried layer is also highly sensitive to parameters such as solvent composition or ambient conditions [28], particle concentration [6], aspect ratio [19] and stiffness [53]. Depending on these factors, deposits can range from uniform coatings or aligned networks to uneven ring stains caused by the coffee-ring effect (CRE), where particles accumulate at the contact line. Uniform deposits are often desirable, especially in coating processes and printing applications [63].
@@
The electrical conductivity of the deposited layer is one of the key properties for applications in printed electronics [9]. Percolation theory plays a key role in understanding and optimizing the conductivity of such morphologies, describing the transition from isolated clusters to a continuous, conductive network, governed by filament concentration, shape, and distribution [8]. Filaments are particularly effective in achieving percolation at low concentrations, as their elongated shapes increase the likelihood of overlap and network formation. Their anisotropic geometry also promotes orientational ordering during deposition, which can enhance electrical, optical, and mechanical properties. This makes them attractive for applications requiring high conductivity and material efficiency [11].
@@
Experimental and theoretical approaches provide valuable insights, but face limitations in isolating and controlling the parameters of drying filament droplets. Numerical simulations overcome these constraints by enabling the systematic variation of individual parameters while keeping others fixed. While spherical particles are well-studied [71, 70, 24], computational modeling of fluid-coupled rod-like structures forming deposition patterns during drying droplets remains largely unexplored [34, 33]. To date, simulations of such systems have often been simplified, for instance by using Monte Carlo (MC) methods that treat filaments as random stick networks without incorporating the fluid dynamics of the evaporating droplet [25, 20].
@@
In this work, we couple the lattice Boltzmann method for fluid dynamics with a bead-spring approach for filaments. Our approach enables the simulation of fluid flow, particle interactions, and network formation, providing insights into filament alignment, clustering, and percolation behavior following solvent drying. We consider the filament suspension in a dilute limit, where self-pinning is negligible, and focus on the stick–slide evaporation mode: droplets evaporate in CCR mode until a critical receding angle $\theta_{re}$ is reached, after which they transition to CCA mode [52, 67]. We investigate how filament length, concentration, stiffness, and evaporation conditions affect percolative network formation through a systematic parameter study. We then analyze the nematic order, radial density profiles, percolation probability, and electrical conductivity of the deposit.
@@ II.1 Color-gradient lattice Boltzmann model
The fluid solver is based on the three-dimensional lattice Boltzmann method with $19$ discrete velocities (D3Q19) [35, 7]. We describe the fluid forming the droplet and the surrounding fluid as independent components. Each component $k\in\{1,2\}$ is described by a distribution function $f_{i}^{k}(\mathbf{x},t)$, which evolves according to the lattice Boltzmann equation,
@@
where $\Omega_{i}^{k}$ is the collision operator, $\mathbf{c}_{i}$ are the discrete velocity vectors, and $i=1,\dots,19$. We set the timestep $\Delta t$ and the lattice constant $\Delta x$ to unity for simplicity. The macroscopic density and velocity of component $k$ are given by
@@
with $\rho_{0}=1$ as the reference density. The unit of mass is defined as $m_{0}=\rho_{0}\,\Delta x^{3}=1$, and the unit of energy as $E_{0}=m_{0}(\Delta x/\Delta t)^{2}=1$.
The CGLB method, enables the simulation of immiscible or partially miscible fluids by modeling interfacial tension and enforcing phase segregation through three substeps: relaxation, perturbation, and recoloring [23, 38, 3]. The total collision operator $\Omega_{i}$ acting on the total distribution function $f_{i}$ is decomposed as
@@
Here, $\tau$ is the dimensionless relaxation time. Through the Chapman–Enskog expansion, $\tau$ is related to the kinematic viscosity $\nu=c_{s}^{2}(\tau-0.5)$, with $c_{s}=\frac{1}{\sqrt{3}}\frac{\Delta x}{\Delta t}$ the speed of sound for the D3Q19 lattice. The equilibrium distribution is given by
@@
where $\rho=\rho_{1}+\rho_{2}$ is the total density and $\mathbf{u}=\frac{1}{\rho}\sum_{i}f_{i}\mathbf{c}_{i}$ is the mixture velocity [35]. The lattice weights $w_{i}$ for the D3Q19 lattice are
@@
Perturbation: Interfacial tension is introduced via a perturbation proportional to the color gradient $\mathbf{G}$ as
@@
where $\phi_{i}$ is the angle between $\mathbf{G}$ and $\mathbf{c}_{i}$. The parameter $A$ sets the surface tension magnitude $\sigma$ according to $A=\frac{9\sigma}{4\tau}$ [38], while the lattice-dependent weights $B_{i}$ are defined as
@@
We note that the unit of surface tension is $E_{0}/\Delta x^{2}=1$ in the simulations.
@@
Solid boundaries are treated using the half-way bounce-back boundary condition, enforcing a no-slip condition with the effective wall located halfway between fluid and solid lattice nodes [35]. To describe wetting on the substrate, we follow the work of Akai et al. [2]. We dynamically adjust the fluid interface normal $\mathbf{n}_{f}$ at the triple line to enforce the contact angle condition. The algorithm computes two candidate normals $\mathbf{v}_{0}$ and $\mathbf{v}_{1}$ based on the geometric relationship between the fluid interface normal $\mathbf{n}_{f}$ and the wall normal $\mathbf{n}_{w}$:
@@ II.2 Evaporation model
Fluid evaporation is implemented following the recent work of Nath et al. [49]. This reaction-limited evaporation model simulates phase change by converting distribution populations at the liquid-vapor interface. This approach is adaptable to any evaporation regime with a known analytical evaporation flux expression and is computationally efficient, as calculations are confined to the interface. Evaporation is modeled by converting resting populations at lattice sites at the interface, identified by a color-gradient magnitude larger than a threshold value $\Gamma=0.305$ [49]:
$f_{0}^{1}(\mathbf{x},t)^{\text{new}}$ $=f_{0}^{1}(\mathbf{x},t)-\varphi\,\Delta t,$ (15)
$f_{0}^{2}(\mathbf{x},t)^{\text{new}}$ $=f_{0}^{2}(\mathbf{x},t)+\varphi\,\Delta t.$
The surface mass flux $\Xi$ across each site is linked to a volumetric sink rate $\varphi=\frac{d\rho}{dt}=S\Xi$, with a correction factor $S:=1/3$ accounting for discretization errors [49]. During each timestep, mass is subtracted from one fluid component and added to another, preserving total mass and momentum.
In printing technologies, evaporation typically occurs in an intermediate regime between the diffusion-limited and reaction-limited limits, which can be adjusted by masking the droplet, varying the relative humidity, the ambient pressure, by imposing an airflow [44] or when molecule transfer through the interface is hindered by the solutes. We restrict our analysis to the dilute regime, assuming negligible particle adsorption at the interface, allowing us to neglect evaporation shielding effects caused by filament accumulation.
In the reaction-limited regime, interfacial kinetics - rather than vapor diffusion - sets the evaporation rate. This commonly occurs on heated substrates or at low pressure [48, 49]. The surface mass flux $\Xi$ follows the Hertz–Knudsen (HK) relation, which provides the kinetic boundary condition linking the interfacial mass flux to the local thermodynamic state of the interface. Using the Clausius-Clapeyron law, the non-dimensional HK relation is given by
$\Xi=\frac{\Lambda\,\rho_{v}\,\mathcal{L}}{T_{\text{sat}}^{3/2}}\sqrt{\frac{M_{v}}{2\pi R_{g}}}\,\left[{\color[rgb]{0,0,0}T_{I}}-T_{\text{sat}}\right],$ (16)
with $\Lambda$ the accommodation coefficient, $\rho_{v}$ the saturation vapor density, $M_{v}$ the molecular mass of the vapor, $R_{g}$ the universal gas constant, $\mathcal{L}$ the latent heat of vaporization, $T_{I}$ the liquid–vapor interface temperature, and $T_{\text{sat}}$ the saturation temperature. In the non-equilibrium one-sided model (NEOS) [37, 48], under the thin-droplet assumption, with negligible substrate thickness and curvature effects, the interface temperature is expressed in terms of the local film height and inserted into the HK relation, yielding the evaporation flux $J$ for thin droplets
@@
In the case of diffusion-limited evaporation of a sessile droplet, vapor transport is assumed quasi-steady and purely diffusive, governed by the Laplace equation [15, 27]
@@
Here, $D_{v}$ is the vapor diffusivity in air and $u(\mathbf{x})$ the vapor concentration field. The boundary conditions are $u=u_{s}$ (saturated vapor concentration) at the droplet surface and $u=u_{\infty}$ (ambient vapor concentration) far away. The droplet is modeled as a spherical cap (small Bond number), and viscous forces are negligible (low capillary number). The resulting evaporation flux $J(r)$ exhibits a divergence near the contact line. Nevertheless, the total evaporation rate remains finite because the singularity is integrable; in addition, real droplets often possess a thin precursor film that regularizes the divergence [58, 51]. For small contact angles $(\theta\to 0)$, the evaporation flux is written as
@@
where $a$ is the droplet base radius and $r$ the radial position (see Fig. 1). For arbitrary angles, an approximate form is [27]
@@
in which $\theta(t)=2\arctan\left(\frac{h_{0}(t)}{a}\right)$ for a spherical droplet in case of semi-steady evaporation. Furthermore, the evaporation flux can be approximated as [48]
@@
In the diffusion-limited regime, the vapor fields of neighboring droplets overlap, leading to a shielding effect that reduces the local evaporation rate. For two identical adjacent droplets the resulting local evaporative flux is expressed as [69]
$J(r,\xi)=J(r)\left[1-\frac{M\sqrt{b^{2}-a^{2}}}{2\pi\left(r^{2}+b^{2}-2rb\cos\xi\right)}\right],$ (22)
where $r$ is the radial coordinate from the center, $\xi$ the azimuthal angle measured from the axis connecting the droplet centers, $a$ is the base radius, and $b$ the center-center-distance between the droplets. The shielding factor $M$ quantifies the reduction of the evaporation flux due to the neighboring droplet and takes the analytical form [69]
@@ II.3 Filament model
The filaments within the droplet are modeled using a bead-spring representation. Each filament is constructed as a sequence of connected beads (monomers) linked by springs. Inter-monomer bonds are governed by the finite extensible nonlinear elastic (FENE) potential, known to accurately model real filament behavior [39]:
@@
Here, $d$ is the separation between adjacent beads, $k_{s}=0.3$ is the spring constant, and $R_{m}=1.2$ is the maximum bond extension. To incorporate excluded volume effects, a WCA-like (Weeks-Chandler-Andersen) potential
@@
is added, where $\varepsilon=0.03$ is the depth of the potential well, and $d_{0}=1$ is the characteristic particle diameter (corresponding to the distance at which the potential vanishes). The stiffness of filaments is modeled using an approximate harmonic potential for second-nearest-neighbor interactions,
@@
Filaments are two-way coupled to the LB fluid via a local momentum-exchange scheme. Each bead experiences a hydrodynamic drag force [1, 29]
@@
with $\gamma=6\pi\eta a_{0}$ representing the friction coefficient for a sphere of radius $a_{0}$ in a fluid of dynamic viscosity $\eta$. Here, $\mathbf{v}_{\text{p}}$ denotes the bead velocity, while $\mathbf{v}_{\text{f}}$ is the fluid velocity interpolated at the bead’s position.
Conversely, the force exerted by the bead on the fluid is mapped back to the lattice using the method proposed by Ahlrichs and Dünweg [1]. This scheme distributes the momentum to the nearest lattice nodes via a regularized delta function. The update rule for the component distributions $f_{i}^{k}$ is given by
@@
The weight function $\delta(\mathbf{x},\mathbf{x}_{\text{p}})$ corresponds to a trilinear interpolation kernel [1], assigning linearly weighted contributions to the nearest lattice nodes in each spatial direction. This formulation locally injects momentum into the fluid while ensuring conservation.
The solvation force, introduced by Sega et al. [59], is used to confine filaments inside the droplet. It depends on the local gradients of the fluid-component densities,
@@
To capture capillary interactions, we calibrate the solvation magnitude $\kappa_{k}$ by balancing the net inward solvation force on the wetted segment $L_{w}$ against the capillary compression. Following the formulation established by Seong et al. [60], we approximate the theoretical capillary force $F_{c}$ as
@@
The prefactor $0.684$ is a geometric scaling factor adopted from [60].
The friction between filaments and the substrate critically influences the final deposition pattern during droplet evaporation. For adhesive surfaces, Derjaguin extended Amontons’ law by including an adhesion-related term [17]. Accordingly, the friction force can be decomposed into external and internal contributions as [21]
@@
where $d$ is the distance from the substrate, and $\varepsilon=1.5\times 10^{-3}$ the potential well depth. This interaction restricts bead mobility, effectively generating internal friction. As evaporation proceeds, frictional forces can exceed solvation forces, immobilizing filaments and producing ring-like deposition patterns [71].